OpenStudy (anonymous):

The ticket office makes $5 profit on each evening ticket sold and $2 on each matinee ticket sold. The ticket office wants to sell at least 60 evening tickets per week and at least 45 matinee tickets per week. Its goal is to earn at least $200 profit in sales per week. Use a graph to show the possible number of tickets sold. Either describe what the graph looks like (include slope, y-intercept, dotted or solid line, and if the shading is above or below the line),

OpenStudy (anonymous):

if you call evening tickets \(x\) and matinee tickets \(y\) then you have a few inequalities \(y\geq 60, x\geq 45\) and also \(5x+2y \geq 200\)